OpenStudy (anonymous):

v(x) = (f(x))^2 Find v''(x).

OpenStudy (anonymous):

I believe that: v'(x) = 2(f(x))*f'(x)

OpenStudy (amistre64):

v = f^2 v' = 2f f' v'' = 2f' f' + 2f f''

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

the (x) parts just clutter the operations